(1 of 12) Match the following: quilibrium constant A. The base formed when an acid deprotonates. product of water B. The ratio of product to reactant concentration of water ionization multiplied by the concentration of water. on Conjugate base C. pH pK idpoint D. The ratio of product to reactant concentration for any given reaction once the equilibrium has been reached
media%2F9ae%2F9ae9ebcf-4c38-4516-9ba6-93
media%2Fb3a%2Fb3afe88c-b35a-4aaf-ba8a-e4
media%2Fd01%2Fd0150651-bde6-47f9-b365-28
Answer #1

1)

Equilibrium Constant :-The ratio of product to reactant  concentration for a given reaction once equilibrium has reached

Ion product of water : The ratio of product to reactant concentration of water ionization multiplied by the concentration of water

Conjugate Base:- The base formed when acid deprotonates

HCl in water produces hydrogen ions (H+)& negatively charged chloride ion (Cl-). The chloride ion (Cl-) is the conjugate base of HCl.

Mid Point : PH= PK ( pK =pH at the midpoint of the titration curve for the acid/base )

2)[H+] = 1*10^ -12 and PH=12

At 298 K ,ionic product of water.= [H+] * [OH-] and by definition At 298 K ,ionic product of water.=10^-14

& in problem given [OH-]= 10^-2 ,So [H+] = 10^-12

Also if [H+] < =10^-6 ,it is acidic ,if [H+] =10^-7 ,it is neutral & if [H+] >10^-7 ,it is alkaline

So here [H+] = 10^-12 ,PH = alkaline and 12 answer B

3)So answer is around 7

4) An organism such as you and I would require a number of buffers all of which need to maintain a pH of ~7.0.